# The Angular Momentum Penrose Inequality

### v2

@@ 1.3 Significance and Relation to Prior Work
Theorem 1.2 establishes the Kerr-sharp angular momentum Penrose inequality—that is, the bound $M_{\mathrm{ADM}}^{2}\geq A/(16\pi)+4\pi J^{2}/A$ with rigidity characterizing the equality case as Kerr—under the hypotheses (H1)–(H4). This result is the first to obtain the exact Kerr Penrose bound without additional horizon integrals or extra assumptions beyond axisymmetry, exterior vacuum, and stable MOTS.
Comparison with prior Penrose-type inequalities involving angular momentum:
Previous works have established Penrose-type inequalities incorporating angular momentum, but with additional terms or under more restrictive assumptions:
  • Khuri–Sokolowsky–Weinstein [67]: Proved a Penrose-type inequality for axisymmetric initial data using harmonic map/Weyl coordinate methods. Their bound takes the form $M_{\mathrm{ADM}}\geq\sqrt{A/(16\pi)+4\pi J^{2}/A}+\mathcal{R}_{\text{rod}}$, where $\mathcal{R}_{\text{rod}}$ is an additional rod integral contribution from the axis structure. This extra term vanishes only when additional conditions on the horizon geometry are imposed. Our result removes this rod integral under the hypotheses (H1)–(H4), yielding the pure Kerr expression.
  • Feng–Yan–Gao–Lau–Yau [25]: Established geometric inequalities for axisymmetric black holes relating mass, angular momentum, and horizon geometry. Their approach uses different techniques (quasi-local mass bounds) and applies under quasi-stationary conditions. Our result applies to general initial data satisfying the dominant energy condition.
  • Alaee–Khuri–Kunduri [3, 4]: Extended Penrose-type inequalities to higher dimensions and different horizon topologies (e.g., $S^{1}\times S^{2}$ black rings). While these works address important generalizations, they do not give the pure Kerr bound in the 4-dimensional setting we consider.
What is genuinely new in this paper:
  • AM-Hawking mass and its monotonicity (Theorems 6.10, 6.22): The functional $m_{H,J}(t)=\sqrt{m_{H}^{2}+4\pi J^{2}/A(t)}$ is new. Its monotonicity combines the standard Hawking mass monotonicity with the Dain–Reiris bound via a “sub-extremality factor.”
  • Angular momentum conservation along AMO flow (Theorem 6.10): While the AMO $p$-harmonic flow is established [1], proving $J(\Sigma_{t})=\mathrm{const}$ along the flow is new and uses co-closedness of the Komar form under vacuum.
  • Axisymmetric Jang equation with twist (Theorem 4.12): We extend the Jang approach to incorporate twist potentials from angular momentum while preserving controlled blow-up behavior on MOTS.
  • Complete rigidity analysis (Theorem 9.1): The synthesis of Mars–Simon tensor methods with foliation rigidity to identify the equality case with Kerr is new.
Relation to foundational prior work:
  • Time-symmetric Penrose inequality (Huisken–Ilmanen [32], Bray [11]): Our result extends theirs to include angular momentum and non-time-symmetric data.
  • Spacetime Penrose inequality (Bray–Khuri [12], Han–Khuri [29]): We build on their Jang equation methods but incorporate the twist perturbation and angular momentum terms.
  • Area-angular momentum inequalities (Dain [20], Dain–Reiris [22]): Their $A\geq 8\pi|J|$ bound is used as an input (not re-derived) to establish sub-extremality control.
  • AMO flows [1]: We use their $p$-harmonic framework but extend it with angular momentum conservation.
@@ 1.4 Organization
The paper is organized as follows:
  • Section 2: Verification that Kerr saturates the inequality
  • Section 3: Overview of the proof strategy
  • Section 4: Axisymmetric Jang equation with twist
  • Section 5: Angular-momentum-modified Lichnerowicz equation
  • Section 6: AMO functional with angular momentum conservation
  • Section 7: Sub-extremality from Dain–Reiris
  • Section 8: Complete proof synthesis
  • Section 9: Rigidity and equality case
  • Section 10: Extensions and open problems
  • Appendix A: Supplementary numerical illustrations
Notation: A comprehensive Glossary of Symbols appears in Section Glossary of Symbols near the end of the paper. For quick reference during a first reading, the principal notation is summarized in Table 1 (page 1).
1.5 Reader’s Guide
For a first reading, we recommend:
  1. 1.
    Read Section 2 to see that Kerr saturates the bound (2 pages).
  2. 2.
    Read Section 3 for the four-stage proof strategy and key diagrams (4 pages).
  3. 3.
    Skim the theorem statements in Sections 47, focusing on the main results (Theorems 4.12, 5.6, 6.10, 6.22, 7.1).
  4. 4.
    Read Section 8 for the complete proof assembly (3 pages).
For verification of technical details, each section contains “Key Estimate Verification Guide” remarks (Remarks 4.20, 5.15, 6.26) that identify the critical estimates and their justifications.
Logical dependencies are summarized in Figure 4. The proof is modular: each of Sections 47 can be read independently given the outputs of previous stages.
Notation help: If you encounter unfamiliar symbols, consult Table 1 below for principal notation and the Glossary of Symbols (Section Glossary of Symbols) for comprehensive definitions.
1.6 Notation Guide
For the reader’s convenience, we collect here the principal notation used throughout the paper.
Symbol Description
Geometric quantities on $(M,g,K)$
$(M^{3},g,K)$ Initial data: Riemannian 3-manifold with metric $g$ and extrinsic curvature $K$
$\eta=\partial_{\phi}$ Axial Killing field generating rotations
$\Gamma$ Rotation axis $\{\eta=0\}$
$\Sigma$ Outermost marginally outer trapped surface (MOTS)
AA Area of $\Sigma$
$J$ Komar angular momentum: $J=\frac{1}{8\pi}\int_{\Sigma}K(\eta,\nu)\,d\sigma$
$\boldsymbol{j}$ Momentum density vector field (boldface)
$\mu$ Energy density: $\mu=\frac{1}{2}(R_{g}+(\mathrm{tr}_{g}K)^{2}-|K|_{g}^{2})$
$M_{\mathrm{ADM}}$ ADM mass at spatial infinity
Jang manifold $(\bar{M},\bar{g})$
$f$ Jang potential (graph function)
$\bar{g}$ Jang metric: $\bar{g}=g+df\otimes df$
$\omega$ Twist 1-form: $\omega_{i}=\epsilon_{ijk}\eta^{j}\nabla^{k}\eta/|\eta|^{2}$
$\tau$ Twist potential (local): $\omega=d\tau$ away from axis
$\mathcal{T}[f]$ Twist perturbation operator in Jang equation
[…]
$\phi$ Conformal factor from AM-Lichnerowicz equation
$\tilde{g}$ Conformal metric: $\tilde{g}=\phi^{4}\bar{g}$
$\Lambda_{J}$ Angular momentum source term: $\Lambda_{J}=\frac{1}{8}|\sigma^{TT}|^{2}$
$R_{\tilde{g}}$ Scalar curvature of $\tilde{g}$ (non-negative by construction)
AMO flow quantities
$u$ $p$-harmonic potential defining the foliation
$\Sigma_{t}$ Level set $\{u=t\}$ for $t\in[0,1]$
$A(t)$ Area of $\Sigma_{t}$
$m_{H}(t)$ Hawking mass: $m_{H}=\sqrt{A/(16\pi)}(1-\frac{1}{16\pi}\int_{\Sigma_{t}}H^{2}\,d\sigma)$
$m_{H,J}(t)$ AM-Hawking mass: $m_{H,J}=\sqrt{m_{H}^{2}+4\pi J^{2}/A}$
Function spaces
$W^{k,p}_{\beta}$ Weighted Sobolev space with exponential weight $e^{\beta t}$ (cylindrical ends)
$C^{k,\alpha}_{-\tau}$ Weighted Hölder space with polynomial decay $r^{-\tau}$ (asymptotically flat ends)
$\lambda_{1}(L_{\Sigma})$ Principal eigenvalue of MOTS stability operator
Table 1: Principal notation used in this paper.
@@ 2 Verification for Kerr Spacetime
For the Kerr spacetime with mass $M$ and spin parameter $a=J/M$ (where $|a|\leq M$ for sub-extremality):
@@
The Kerr initial data $(g_{\mathrm{Kerr}},K_{\mathrm{Kerr}})$ on a Boyer–Lindquist constant-$t$ slice belongs to the weighted spaces required by Theorem 1.2. Specifically, in the coordinates $(r,\theta,\phi)$ with $r>r_{+}$ (exterior region):
@@
The regularity extends across the bifurcation sphere (MOTS) by standard analysis of the Kerr metric in horizon-penetrating coordinates (e.g., Kerr–Schild). Thus Kerr data satisfies all hypotheses of Theorem 1.2 for $0<|a|<M$ (strictly sub-extremal) and satisfies hypotheses (H1)–(H3) for all $|a|\leq M$.
@@
3 Proof Strategy: Overview
The proof uses the four-stage Jang–conformal–AMO method, extending techniques from the spacetime Penrose inequality literature [12, 29, 1].
3.1 Proof Roadmap
For readers seeking a quick overview of the logical structure, the proof follows this dependency chain:
Proof Roadmap: From Hypotheses to Conclusion $\boxed{\text{(H1) DEC}}\xrightarrow{\text{Jang eq.}}R_{\bar{g}}\geq 0\xrightarrow{\text{AM-Lich.}}\phi>0\text{ exists}\xrightarrow{\text{conformal}}R_{\tilde{g}}\geq 0$ $\boxed{\text{(H2) Axisym.}}+\boxed{\text{(H3) Vacuum}}\xrightarrow{d^{\dagger}\alpha_{J}=0}J(t)=J\text{ (conserved)}$ $\boxed{\text{(H4) Stable MOTS}}\xrightarrow{\text{Dain--Reiris}}A(0)\geq 8\pi|J|\xrightarrow{A^{\prime}(t)\geq 0}A(t)\geq 8\pi|J|\;\forall t$ $R_{\tilde{g}}\geq 0+J\text{ conserved}+\text{sub-extremal}\xrightarrow{\text{AMO}}\frac{d}{dt}m_{H,J}(t)\geq 0$ $m_{H,J}(1)=M_{\text{ADM}}\geq m_{H,J}(0)=\sqrt{\frac{A}{16\pi}+\frac{4\pi J^{2}}{A}}\quad\checkmark$
3.2 Comparison with Prior Penrose Inequality Proofs
The following table compares our approach with the two established proofs of the (non-rotating) Riemannian Penrose inequality:
Feature Huisken–Ilmanen [32] Bray [11] This Paper
Flow type Inverse mean curvature (IMCF) Conformal flow $p$-harmonic (AMO)
Handles $J\neq 0$? No No Yes
Data type Time-symmetric ($K=0$) Time-symmetric ($K=0$) General $(g,K)$
Symmetry required? None None Axisymmetry
Matter assumption $R_{g}\geq 0$ $R_{g}\geq 0$ DEC + vacuum exterior
Boundary condition Weak solution jumps Horizons shrink to points Cylindrical ends
Monotonic quantity Hawking mass $m_{H}$ Isoperimetric mass AM-Hawking mass $m_{H,J}$
Rigidity case Schwarzschild Schwarzschild Kerr
Multiple horizons? Yes (jumps) Yes (conformal) One (outermost)
Regularity required Weak solutions $C^{2}$ $C^{2,\alpha}$ weighted
Table 2: Comparison of Penrose inequality proof methods. The key advantage of our approach is the ability to handle rotating black holes ($J\neq 0$) with general initial data, at the cost of requiring axisymmetry and vacuum exterior hypotheses.
Remark 3.1 (Self-Contained Proof).
The proof is self-contained in that it does not require prior results about the Penrose inequality as inputs. Each stage uses established techniques from geometric analysis: Han–Khuri [29, Theorem 1.1, Proposition 4.5] for Jang existence, standard elliptic theory for the Lichnerowicz equation, AMO [1, Theorem 1.1] for monotonicity, and Dain–Reiris [22, Theorem 1] for the area-angular momentum inequality on MOTS. The principal novelty lies in the synthesis of these methods and the introduction of the AM-Hawking mass for the rotating case.
3.3 The Four Stages
The proof proceeds through four main stages, each building on the previous. We summarize the construction before presenting the technical details.
Stage 1: Jang Equation Axisymmetric with twist $\to$ Jang manifold $(\bar{M},\bar{g})$ Stage 2: Conformal Sealing AM-Lichnerowicz equation $\to$ $\tilde{g}=\phi^{4}\bar{g}$ with $R_{\tilde{g}}\geq 0$ Stage 3: AMO Flow $J$-modified Hawking mass $\to$ Monotonicity of $m_{H,J}(t)$ Stage 4: Boundary Values $m_{H,J}(0)\geq\sqrt{A/16\pi+4\pi J^{2}/A}$ $m_{H,J}(1)=M_{\mathrm{ADM}}$
Figure 1: The four-stage Jang–conformal–AMO proof strategy. Each stage transforms the geometric data while preserving or establishing key properties needed for the inequality.
@@ 3.3 The Four Stages
Figure 2: Schematic showing how the AM-Penrose inequality emerges from the geometric constructions.
[…]
Figure 3: The chain of manifold transformations from physical initial data to the conformal manifold with non-negative scalar curvature.
[…]
Figure 4: Logical dependencies among the key results. Each arrow indicates how one result is used to derive the next.
3.4 Key Modifications from Spacetime Penrose Proof
Component Standard Penrose AM-Penrose
Jang equation $H_{\Gamma}=\mathrm{tr}_{\Gamma}K$ Add twist source $S_{\omega}[f]$
Lichnerowicz $8\Delta\phi+R\phi=0$ Add $\Lambda_{J}\phi^{-7}$ term
Monotonic functional Hawking mass $m_{H}$ AM-Hawking mass $m_{H,J}$
Conservation Area monotonicity Area mono. + $J$ conservation
Boundary at $\infty$ $m_{H}(1)=M_{\mathrm{ADM}}$ $m_{H,J}(1)=M_{\mathrm{ADM}}$
Table 3: Comparison of proof components between the standard Penrose inequality (for non-rotating black holes) and the angular momentum Penrose inequality (rotating case). Each row shows how a key ingredient is modified to incorporate angular momentum $J$.
3.5 Four Technical Theorems
The proof requires establishing four technical results:
@@ 3.5 Four Technical Theorems
    Jang Existence4): The axisymmetric Jang equation with twist as a lower-order perturbation admits a solution with cylindrical ends at the MOTS, preserving angular momentum information.
@@
    AM-Lichnerowicz5): The angular-momentum-modified Lichnerowicz equation has a unique positive solution $\phi$ with $\phi|_{\Sigma}=1$ and $\phi\to 1$ at infinity, yielding a conformal metric with $R_{\tilde{g}}\geq 0$.
@@
    $J$ Conservation6): For axisymmetric vacuum data, the Komar angular momentum $J(t)=J$ is constant along the AMO flow (by Stokes’ theorem applied to the co-closed Komar form).
@@
    Sub-Extremality7): The Dain–Reiris inequality [22] gives $A(t)\geq 8\pi|J|$ for all $t$, ensuring the sub-extremality factor in the monotonicity formula is non-negative.
3.6 Key Estimates Summary
For readers verifying this proof, we provide a summary of the critical estimates and their locations:
Estimate Statement Location
Twist perturbation bound $|\mathcal{T}|=O(s)$ as $s\to 0$ near MOTS Thm 4.12, Step 2c
Jang blow-up rate $f(s,y)=C_{0}\ln s^{-1}+O(1)$, $C_{0}=|\theta^{-}|/2$ Thm 4.12(ii)
Indicial root positivity $\lambda_{0}(-8\Delta_{\Sigma}+R_{\Sigma})>0$ Lem 5.5, Step 3
Conformal factor decay $|\phi-1|=O(e^{-\kappa t})$ on cylindrical end Lem 5.12, Step (ii)
Flux vanishing $\lim_{R\to\infty}\int_{S_{R}}\phi^{2}\partial_{\nu}\phi\,d\sigma\geq 0$ Lem 5.11
Co-closedness of Komar form $d^{\dagger}\alpha_{J}=\boldsymbol{j}\cdot\eta=0$ (vacuum) Thm 6.10, Step 5
Sub-extremality factor $(1-(8\pi|J|/A)^{2})\geq 0$ when $A\geq 8\pi|J|$ Thm 6.22, Step 8g
Global AM-Hawking bound $M_{\mathrm{ADM}}\geq m_{H,J}(t)\geq m_{H,J}(0)$ via comparison Rem 6.23, Parts I–IV
$p$-harmonic uniform bounds $\|u_{p}\|_{C^{1,\alpha}(K)}\leq C(K)$ uniformly in $p\in(1,2]$ Lem 6.28
Table 4: Critical estimates and their locations in the proof. These bounds are essential for verifying the main theorem; each estimate is used in the subsequent stages of the argument. The “Location” column provides precise references to where each estimate is established.
@@ 4 Stage 1: Axisymmetric Jang Equation
Remark 4.1 (Bounded Geometry at the Axis).
Throughout this section, we assume that the initial data has bounded geometry at the symmetry axis $\Gamma=\{\eta=0\}$. Explicitly: the metric $g$ and extrinsic curvature $K$ extend smoothly to the axis, and in Cartesian coordinates $(x,y,z)$ with $\rho=\sqrt{x^{2}+y^{2}}$, all metric coefficients and their derivatives are bounded functions of $(x,y,z)$. This implies that the twist 1-form $\omega$ satisfies $|\omega|_{g}=O(1)$ and $|\nabla\omega|_{g}=O(1)$ as $\rho\to 0$—see (18) and the axis regularity conditions (AR1)–(AR3) below. This regularity is automatic for data arising from smooth stationary axisymmetric spacetimes (e.g., Kerr).
@@ 4.1 Function Spaces and Regularity Framework
For $k\in\mathbb{N}_{0}$, $\alpha\in(0,1)$, and weight $\tau\in\mathbb{R}$, define the weighted Hölder space on an asymptotically flat manifold $(M,g)$ with asymptotic radial coordinate $r(x):=|x|$ in the end:
$C^{k,\alpha}_{-\tau}(M):=\{u\in C^{k,\alpha}_{\mathrm{loc}}(M):\|u\|_{C^{k,\alpha}_{-\tau}}<\infty\},$
@@
$\|u\|_{C^{k,\alpha}_{-\tau}}:=\sum_{|\beta|\leq k}\sup_{x\in M}\langle r(x)\rangle^{\tau+|\beta|}|D^{\beta}u(x)|+[D^{k}u]_{\alpha,-\tau-k-\alpha},$
@@
$v]_{\alpha,\delta}:=\sup_{\begin{subarray}{c}x\neq y\\ d(x,y)<\mathrm{inj}(M)/2\end{subarray}}\min(\langle r(x)\rangle,\langle r(y)\rangle)^{-\delta}\frac{|v(x)-v(y)|}{d(x,y)^{\alpha}}.$
Here $\mathrm{inj}(M)$ denotes the injectivity radius. A function $u\in C^{k,\alpha}_{-\tau}(M)$ satisfies $|u(x)|=O(r^{-\tau})$ as $r\to\infty$.
@@
    Metric decay: $g_{ij}-\delta_{ij}\in C^{2,\alpha}_{-\tau}(M\setminus K_{0})$, i.e.,
@@
    Extrinsic curvature decay: $K_{ij}\in C^{1,\alpha}_{-\tau-1}(M\setminus K_{0})$, i.e.,
@@
The condition $\tau>1/2$ ensures convergence of the ADM integral: the integrand is $O(R^{-\tau-1})$, so the surface integral is $O(R^{2-\tau-1})=O(R^{1-\tau})\to 0$ as $R\to\infty$ when $\tau>1$; the weaker condition $\tau>1/2$ suffices by more refined analysis using the constraint equations (see [10, Theorem 4.2]).

### v3

@@ 1.3 Significance and Relation to Prior Work
Theorem 1.2 establishes a geometric inequality incorporating both horizon area and angular momentum for axisymmetric vacuum initial data. We describe the contributions of this paper and their relationship to prior work, with appropriate context.
1.4 Relation to Prior Work
The Angular Momentum Penrose Inequality has a substantial history dating to Penrose’s original proposal [48]. Our work synthesizes existing methods—specifically the Jang equation approach of Bray–Khuri [11] and Han–Khuri [29], the area-angular momentum inequalities of Dain–Reiris [22], and the $p$-harmonic flow of Agostiniani–Mazzieri–Oronzio [1]—into a complete proof for the axisymmetric vacuum case.
Technical contributions:
  • AM-Hawking mass functional: We introduce the functional $m_{H,J}(t)=\sqrt{m_{H}^{2}+4\pi J^{2}/A(t)}$ and establish its monotonicity by combining Hawking mass monotonicity with the Dain–Reiris sub-extremality bound.
  • Angular momentum conservation: We prove that $J(\Sigma_{t})$ is constant along the $p$-harmonic level set flow under the vacuum hypothesis (Theorem 6.14).
  • Axisymmetric Jang equation: We extend the Han–Khuri analysis to incorporate twist potentials, establishing the necessary decay estimates for the twist perturbation near the MOTS (Theorem 4.12).
Relation to prior work:
  • Time-symmetric Penrose inequality (Huisken–Ilmanen [32], Bray [10]): These foundational works established the $J=0$ case using IMCF and conformal flow respectively. Our result extends to $J\neq 0$ using different flow methods (AMO $p$-harmonic flow).
  • Spacetime Penrose inequality (Bray–Khuri [11], Han–Khuri [29]): The Jang equation approach and the scalar curvature identity $R_{\bar{g}}\geq 2|\boldsymbol{j}|^{2}$ are due to Bray–Khuri. Han–Khuri developed the MOTS blow-up analysis that we rely upon.
  • Area-angular momentum inequalities (Dain [20], Dain–Reiris [22]): The fundamental bound $A\geq 8\pi|J|$ for stable MOTS is their result. We use this as an input—we do not reprove it. Our contribution is showing how this bound integrates into the mass inequality proof.
  • AMO flows [1]: The $p$-harmonic flow and Hawking mass monotonicity are established by Agostiniani–Mazzieri–Oronzio. We adapt their framework to track angular momentum.
  • Mars–Simon uniqueness [83, 84]: The Simon tensor characterization of Kerr is classical. We apply these methods in our specific geometric setting.
@@ 1.5 Organization
The paper is organized as follows. In Section 2, we verify the inequality for the Kerr spacetime. Section 3 outlines the proof strategy. Section 4 establishes the existence of solutions to the axisymmetric Jang equation. Section 5 treats the angular-momentum-modified Lichnerowicz equation. Section 6 details the AMO flow and angular momentum conservation. Section 7 incorporates the sub-extremality bound. Section 8 synthesizes these results to prove the main theorem. Finally, Section 9 addresses the rigidity case, and Section 10 discusses extensions.
@@ 2 Verification for Kerr Spacetime
We work in geometric units ($G=c=1$). For the Kerr spacetime with mass $M$ and spin parameter $a=J/M$ (where $|a|\leq M$ for sub-extremality):
@@
The Kerr initial data $(g_{\mathrm{Kerr}},K_{\mathrm{Kerr}})$ on a Boyer–Lindquist constant-$t$ slice is asymptotically flat and satisfies the standard Kerr falloff. In particular, in the coordinates $(r,\theta,\phi)$ with $r>r_{+}$ (exterior region):
@@
Moreover, by working in horizon-penetrating coordinates (e.g., Kerr–Schild or ingoing Eddington–Finkelstein), one can extend the induced geometry across the horizon (and in particular across the bifurcation sphere) in a smooth fashion; see, for example, standard GR references such as [17].
Thus Kerr data satisfies the hypotheses of Theorem 1.2 in the strictly sub-extremal case $0<|a|<M$. In the extremal limit $|a|=M$, one should treat near-horizon regularity and the precise formulation of “slice through the horizon” with additional care; for this reason we only claim (H1)–(H3) uniformly for the full range $|a|\leq M$.
@@
Remark 2.7 (Summary: Kerr Saturation Verified).
Substituting the Kerr horizon parameters ($A=8\pi M(M+\sqrt{M^{2}-a^{2}})$ and $J=aM$) into the inequality yields exactly $M_{\mathrm{ADM}}=M$, confirming the bound is sharp. This verification is a necessary consistency check: if Kerr failed to saturate the bound, the conjecture would be false. The explicit algebraic identity $A/(16\pi)+4\pi J^{2}/A=M^{2}$ for Kerr (Theorem 2.3) determines the correct functional form of the angular momentum Penrose inequality.
3 Proof Strategy
The proof employs a four-stage method extending the techniques of Bray–Khuri [11], Han–Khuri [29], and Agostiniani–Mazzieri–Oronzio [1]. The strategy transforms the initial data $(M,g,K)$ into a manifold with non-negative scalar curvature, on which a monotonicity argument can be applied.
3.1 Overview of the Construction
The proof proceeds through the following stages:
  1. 1.
    Jang Equation: We solve the axisymmetric Jang equation to construct a graph $\Sigma$ in the product manifold $M\times\mathbb{R}$. The resulting Jang manifold $(\bar{M},\bar{g})$ has a cylindrical end corresponding to the original MOTS and satisfies a scalar curvature inequality involving the energy density.
  2. 2.
    Conformal Deformation: We solve a generalized Lichnerowicz equation for a conformal factor $\phi$, producing a metric $\tilde{g}=\phi^{4}\bar{g}$ with non-negative scalar curvature $R_{\tilde{g}}\geq 0$. The source term in this equation involves the angular momentum, ensuring that the geometry encodes the rotation.
  3. 3.
@@ 3.1 Overview of the Construction
  4. 4.
    Monotonicity and Boundary Values: We define the AM-Hawking mass $m_{H,J}(t)$ and prove its monotonicity along the flow. The sub-extremality condition $A\geq 8\pi|J|$ ensures the non-negativity of the derivative. Evaluating the mass at the horizon ($t=0$) and at infinity ($t=1$) yields the inequality.
3.2 Comparison with Previous Methods
Our approach differs from the standard Riemannian Penrose inequality proofs [32, 10] by incorporating angular momentum into the flow and the curvature condition. Unlike the inverse mean curvature flow or conformal flow used in the time-symmetric case, we use the $p$-harmonic flow, which allows for a natural coupling with the angular momentum conservation via the level set geometry.
3.3 Technical Components
The proof relies on four main technical results, which we establish in the subsequent sections:
@@ 3.3 Technical Components
    Jang Existence4): The existence of solutions to the axisymmetric Jang equation with controlled blow-up at the MOTS.
@@
    AM-Lichnerowicz5): The solvability of the conformal problem with the angular momentum source term.
@@
    Conservation6): The conservation of $J$ along the flow.
@@
    Sub-Extremality7): The preservation of the sub-extremality condition $A(t)\geq 8\pi|J|$ along the flow.
@@ 4.1 Function Spaces and Regularity Framework
For $k\in\mathbb{N}_{0}$, $\beta\in(0,1)$, and weight $\tau\in\mathbb{R}$, define the weighted Hölder space on an asymptotically flat manifold $(M,g)$ with asymptotic radial coordinate $r(x):=|x|$ in the end:
$C^{k,\beta}_{-\tau}(M):=\{u\in C^{k,\beta}_{\mathrm{loc}}(M):\|u\|_{C^{k,\beta}_{-\tau}}<\infty\},$
@@
$\|u\|_{C^{k,\beta}_{-\tau}}:=\sum_{|\beta|\leq k}\sup_{x\in M}\langle r(x)\rangle^{\tau+|\beta|}|D^{\beta}u(x)|+[D^{k}u]_{\beta,-\tau-k-\beta},$
@@
$v]_{\beta,\delta}:=\sup_{\begin{subarray}{c}x\neq y\\ d(x,y)<\mathrm{inj}(M)/2\end{subarray}}\min(\langle r(x)\rangle,\langle r(y)\rangle)^{-\delta}\frac{|v(x)-v(y)|}{d(x,y)^{\beta}}.$
Here $\mathrm{inj}(M)$ denotes the injectivity radius. A function $u\in C^{k,\beta}_{-\tau}(M)$ satisfies $|u(x)|=O(r^{-\tau})$ as $r\to\infty$.
@@
    Metric decay: $g_{ij}-\delta_{ij}\in C^{2,\beta}_{-\tau}(M\setminus K_{0})$, i.e.,
@@
    Extrinsic curvature decay: $K_{ij}\in C^{1,\beta}_{-\tau-1}(M\setminus K_{0})$, i.e.,
@@
The condition $\tau>1/2$ ensures convergence of the ADM integral: the integrand is $O(R^{-\tau-1})$, so the surface integral is $O(R^{2-\tau-1})=O(R^{1-\tau})\to 0$ as $R\to\infty$ when $\tau>1$; the weaker condition $\tau>1/2$ suffices by more refined analysis using the constraint equations (see [9, Theorem 4.2]).